Customers using a self-service soda dispenser take an average of 12 ounces of soda with an SD of 2 ounces. Assume that the amount would be normally distributed.

What is the probability that a randomly selected customer takes over 10 ounces of soda?
What is the probability that a randomly selected customer takes between 13 to 14 ounces of soda?
How many of the next 100 customers will take an average of less than 12.24 ounces?
1
Expert's answer
2019-11-11T13:25:21-0500

Let "X=" the number of ounces of soda: "X\sim N(\mu, \sigma^2)"

Then


"Z={X-\mu \over \sigma}\sim N(0,1)"

Given that "\mu=12 \ ounces , \sigma=2\ ounces"


"P(X>10)=1-P(X\leq 10)=1-P(Z\leq {10-12 \over 2})="

"=1-P(Z\leq -1)\approx1-0.158655\approx0.8413"

The probability that a randomly selected customer takes over 10 ounces of soda is "0.8413"


"P(13<X<14)=P(X<14)-P(X<13)="


"=P(Z< {14-12 \over 2})-P(Z< {13-12 \over 2})="

"=P(Z< 1)-P(Z<0.5)\approx0.841348-0.691462\approx"

"\approx0.1499"

The probability that a randomly selected customer takes between 13 to 14 ounces of soda is "0.1499"


Given that "n=100"


"P(X<12.24)=P(Z<{12.24-12 \over 2/\sqrt{100}})=P(Z<1.2)\approx"

"\approx0.8849"

"0.8849\cdot100=88.49\approx88"

88 of the next 100 customers will take an average of less than 12.24 ounces.